write an equation in slope intercept form of the line that passes through the point and is perpendicular to the graph and given equation .
Through (-2,-4) perpendicular to x-2y=4

Answer by mananth(16946)

x-2y = 4
Find the slope of this line

-2y=-x+ 4
Divide by -2
y = 1/ 2 x-2
Compare this equation with y=mx+b
slope m = 1/ 2

The slope of a line perpendicular to the above line will be the negative reciprocal -2
m1*m2=-1


m= -2 ,point ( -2 , -4 )
Find b by plugging the values of m & the point in
y=mx+b
-4=4+b
b= -8
m= -2
The required equation is y=-2x-8
m.ananth@hotmail.ca